OpenStudy (anonymous):

The function ƒ(x) = −(x + 3)2 − 4 is not one-to-one. Find a portion of the domain where the function is one-to-one and find an inverse function. The restricted domain for ƒ is: a. [-3-,∞) and F^-1(x)=-3-√x+4 b.[-3-,∞) and F^-1(x)=-3+√x+4 c.[-3-,∞) and F^-1(x)=-3-√-x-4 d.[-3-,∞) and F^-1(x)=-3+√-x-4

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hello

OpenStudy (misty1212):

all of your domains are the same, so no worries there it is only a matter of finding the inverse right?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

ok lets solve \[x=-(y+3)^2-4\] for \(y\)

OpenStudy (misty1212):

add four get \[x+4=-(y-3)^2\]

OpenStudy (misty1212):

change the sign of both sides get \[-x-4=(y-3)^2\]

OpenStudy (misty1212):

take the square root of both sides get \[\sqrt{-x-4}=y-3\] normally it would be \(\pm\sqrt{-x-4}\) but we have restricted the domain so only the plus

OpenStudy (anonymous):

oh ok I get it

OpenStudy (misty1212):

i made a typo

OpenStudy (misty1212):

\[\sqrt{-x-4}=y+3\]

OpenStudy (anonymous):

thats how the 3 becomes negative

OpenStudy (misty1212):

yeah but let me check something else

OpenStudy (anonymous):

ok